# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

### v1

@@ 7 Conclusion
The most important insight that we conclude for the paper is, preprocessing SGD might already be enough for LLM training tasks. This presents an elegant solution to address the memory-performance dilemma of training LLMs using adaptive optimizers like Adam. The other insight comes from the fact that the key components of the SWAN optimizer were designed using key insights from LLM dynamics under a simplified setting. In that sense, this paper serves as a proof-of-concept on the value of studying the dynamics of LLM architectures when designing optimizers. We hope both research insights can inspire more research alone this line to further advance this field.
@@ Appendix A Desired properties of adaptive optimizers
Appendix B Theoretical Analysis
B.1 Analyzing the GradNorm: A Transformer Learning Dynamics Perspective
We take the transformer LLM architecture as an example, and demonstrate that $\mathtt{GradNorm}$ effectively removes the time-variant components of gradient noises. For our analysis, we assume the following simplified transformer block (STB) architecture recently proposed in Tian et al. (2023), which contains the major building blocks of transformers and relies on relatively less restrictive assumptions and simplifications, making them suitable for theoretical analysis for transformer dynamics. It is defined as follows:
Definition 1 (Simplified Transformer Block (STB)).
Given the input activation ${\bm{x}}\in\mathbb{R}^{M_{C}\times 1}$, query token index $q$, context embedding matrix ${\mathbf{U}}_{C}\in\mathbb{R}^{d\times M_{C}}$, and the query embedding ${\bm{u}}_{q}\in\mathbb{R}^{d\times 1}$, the simplified transformer block computes the output ${\bm{h}}\in\mathbb{R}^{n\times 1}$ as follows:
${\bm{h}}=\phi\left({\mathbf{W}}^{\top}\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)\right),$
where the attention logits ${\bm{z}}_{q}\in\mathbb{R}^{M_{C}\times 1}$ are given by
$z_{ql}={\bm{u}}_{q}^{\top}{\mathbf{W}}_{Q}^{\top}{\mathbf{W}}_{K}u_{l},$
with ${\mathbf{W}}_{Q},{\mathbf{W}}_{K}\in\mathbb{R}^{d\times d}$ being weight matrices for the queries and keys, respectively, ${\mathbf{W}}\in\mathbb{R}^{d\times n}$ is the weight matrix for the feedforward network, and $\phi$ is a nonlinearity function such as ReLU.
Given a STB, we consider the following standard mini-batch learning dynamics. Define the conditional expectation $\mathbb{E}_{q=m}[\cdot]:=\mathbb{E}[\cdot|q=m]$. Consider the dynamics of the weight matrix $W$ and the attention logits $z_{q}$, if we train the model with a batch of inputs that always end up with query $q[i]=m$. The weight update for $W$ and $z_{q}$ are given by the following noisy updates:
$\dot{{\mathbf{W}}}^{(t)}=\mathbb{E}_{q=m}\left[{\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}\right]+{\bm{\varepsilon}}^{(t)},\quad\dot{{\bm{z}}}_{m}^{(t)}=\mathbb{E}_{q=m}\left[\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}\right]+{\bm{\varsigma}}^{(t)},$ (10)
Where ${\bm{f}}^{(t)}=\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q}^{(t)})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)$, $({\bm{h}}^{(t)})^{\prime}=\phi^{\prime}(({\mathbf{W}}^{(t)})^{\top}{\bm{f}}^{(t)})$ is the derivative of the current activation, ${\mathbf{G}}_{{\bm{h}}}^{(t)}=\nabla_{{\bm{h}}^{(t)}}\mathcal{L}$ is the gradient of the loss function $\mathcal{L}$ with respect to the hidden activation ${\bm{h}}^{(t)}$, and ${\bm{g}}_{{\bm{f}}^{(t)}}^{(t)}=\sum_{k}{\bm{g}}_{{\bm{h}}_{k}^{(t)}}^{(t)}({\bm{h}}^{(t)}_{k})^{\prime}{\bm{w}}_{k}^{(t)}$ is the sum of the gradients with respect to the attention logits. Here, ${\bm{w}}_{k}^{(t)}$ is the $k$-th column of ${\mathbf{W}}^{(t)}$, ${\bm{g}}_{h_{k}^{(t)}}^{(t)}[i]$ be the backpropagated gradient sent to node $k$ at sample $i$. Finally, ${\bm{\varepsilon}}$ and ${\bm{\varsigma}}$ are random noises that are induced by the minibatch sampling.
Remark
Given STB under the dynamics of Equation 10, the distribution of noises ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ generally vary across time, since both the integrand within expectation, ${\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}$ and $\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}$ vary across time. However, the following result on $\mathtt{GradNorm}$ suggests that this time variability of ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ are completely removed by $\mathtt{GradNorm}$:
Theorem 1 ($\mathtt{GradNorm}$ stabilizes gradient distribution across time for STB).
Consider the STB (Definition 1). Assuming we inherit the assumptions in Theorem 1 of Tian et al. (2023), as described in Appendix B. Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole. Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}(\top{\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
In other words, the covariance structure of $\tilde{{{\mathbf{G}}}}$ is identical across all time steps $t$, achieving distributional stability across time. The same relationship also holds for the gradient of attention score $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}(\top{\bm{x}}^{(t)})}{\partial{\bm{z}}_{q}})$.
Theorem 1 suggests that $\mathtt{GradNorm}$ makes use of the inductive bias induced by transformer architectures and directly removes the time-variant components in gradient covariance structures caused by mini-batch sampling. In practice, $\mathtt{GradNorm}$ makes the following approximation to Theorem 1:For MLP layer $\mathtt{GradNorm}$ is applied on ${\mathbf{W}}$ instead of ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$ as a whole, since such grouping does not exist in practical transformer architectures. However, it still works reasonably well in practice (Section 5.1, Section 5.6).
B.2 Analyzing GradWhitening Pt. I: A Transformer Learning Dynamics Perspective
In this section, we take the perspective of transformer learning dynamics, and show that how $\mathtt{GradWhitening}$ can be naturally derived from the simplified transformer model defined in Definition 1.
Let us first consider the dynamics of the Hessian matrix of STB. We will show that: i), it turns out that at equilibirum, the Hessian is highly structured; ii), under some conditions, $\mathtt{GradWhitening}$ recovers Newton’s method-like behaviors under such Hessian structures. Below we first present the following result on the Hessian structure of the simplified transformer block at equilibrium:
Consider a STB (1), trained by full-batch dynamics (i.e., removing noise in Equation 10). Assume we inherit all the assumptions from Theorem 1 of Tian et al. (2023). Consider the following Hessian matrix:
${\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}=\frac{\partial\mathcal{L}}{\partial{w_{sk}}\partial{w_{s^{\prime}k^{\prime}}}}$
where $1\leq s,s^{\prime}\leq m$, $1\leq k,k^{\prime}\leq n$. Then, as $t\xrightarrow[]{}\infty$, we have the following shared Hessian structure along the diagonal blocks:
Proposition 1 shows that, under a simplified setting of transformer, because of the winner-takes-all behavior in the attention model, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, after normalization. This result is verified in our numerical experiment as shown in Figure 9.
Given by the inductive bias that equality across diagonal blocks of Hessians at transformer learning equilibrium, in the next theorem we show that $\mathtt{GradWhitening}$ can be derived as a second-order method under this structural assumption
Theorem 2 ($\mathtt{GradWhitening}$ as structural approximation of Hessian).
Consider the optimization problem given by $\min_{W\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ where ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, with $m<n$. Assume:
  • The local Hessian $H$ of $f$ has shared block-diagonal, such that ${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
  • ${\mathbf{W}}^{(t)}$ is initialized as a orthogonal matrix.
Then (under 2nd order approximation) $\mathtt{GradWhitening}$’s modified gradient descent (Equation 13) on $f$ recovers the update of Newton’s method, without explicitly estimate second-order quantities.
To summarize, $\mathtt{GradWhitening}$ recovers a Newton’s method-like behavior under the highly structured Hessian regime (i.e., there exists $\tilde{{\mathbf{H}}}$ such that ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$), which is partially satisfied in transformer equilibrium dynamics (see remark below). Compared to directly applying accelerated second-order methods such as the Gauss-Newton method ($\mathcal{O}(Bm^{2}n^{2}+m^{3}n^{3})$, $B$ being batch size) or conjugate gradient method ($\mathcal{O}(BTmn)$ complexity, $T=$ 10 $\sim$ 100, with worst case complexity $\mathcal{O}(Bm^{2}n^{2})$), $\mathtt{GradWhitening}$ should be much faster ($\mathcal{O}(mn\min\{m,n\})$), especially under the Newton-Schulz iteration.
Remark: limitations
Proposition 1 deviates from the strict assumptions of Theorem 2 in that: Proposition 1 only asserts that diagonal direction of Hessian shares the identical structure, but does not assert that the off-diagonal blocks are zero, without further assumptions on $g_{h_{k}}$. However, given that it is standard practice to ignore off-diagonal elements in many practical optimizers including Adam, we hereby argue that ignoring off-diagonal blocks is already a better approximation.
B.3 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
@@ B.1 Analyzing the GradNorm: A Transformer Learning Dynamics Perspective
Unlike the convergence rates of GD and Adam presented in Zhang et al. (2024), as well as Theorem 4 and Corollary 1 in Appendix, the optimal convergence rate (15) of $\mathtt{GradWhitening}$ no longer explicitly depends on the condition number $\kappa$ of $H$. In fact, consider a lower bound $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$, since trace of $H$ appear both in the nominator and denominator, we expect that to be more robust to ill-conditioned problems. For example, consider the specific initialization ${\mathbf{W}}_{\text{whitened}}^{(t)}=cI$, it is straightforward to show that $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\perp\kappa$, which is completely disentangled from the condition number. Hence $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$ would not shrink as $\kappa\xrightarrow{}\infty$. See Proposition 2 for less extreme situations.
@@
As pointed out by Zhang et al. (2024) and Da Silva & Gazeau (2020), Adam with $\beta_{2}<1$ will have issues with convergence, which will not be completely removed even with lr decay. Therefore, we will not discuss the case of $\beta_{2}<1$ to avoid the complication.
[…]
@@ Appendix D Proof of Theorem 3
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}\mathcal{P}({\mathbf{H}}{\mathbf{O}}),$
noticing that $\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}=1$, and $\mathcal{P}({\mathbf{H}}{\mathbf{O}})=\mathcal{P}({\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{O}})={\mathbf{Q}}{\mathbf{Q}}^{\top}{\mathbf{O}}={\mathbf{O}}$. Hence:
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-{\mathbf{O}}=\mathbf{0}={\mathbf{W}}^{*}.$
@@ Appendix F Proof of Proposition 3
Theorem 4 (Contraction factor lower bound for gradient descent, generalized based on Zhang et al. (2024)).
@@
The proposition 1 in Zhang et al. (2024) has shown that the lower bound holds for diagonal positive definite Hessian ${\mathbf{H}}$. To show that the lower bound holds for a general positive definite Hessian ${\mathbf{H}}$ we will reformulate the problem to align with the setup in diagonal case (Prposition 1 of Zhang et al. (2024)).
@@
Thus, by applying Proposition 1 of Zhang et al. (2024) to this transformed problem, we conclude that there exists initial point such that the lower bound on the contraction rate
@@
The update rule of Adam with $\beta_{1}=0$ and $\beta_{2}=1$ is given by Zhang et al. (2024):

### v2

@@ 7 Conclusion
We introduced SWAN, a stateless optimizer for LLM training that combines two well-known operators applied to raw gradients, $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$, to stabilize the stochasticity of gradient distribution and neutralizing the local geometry of loss landscape, respectively. Through theoretical analysis and empirical evidence, we showed that SWAN reduces memory usage while achieving on-par or even better performance compared to Adam on LLaMA pre-training tasks. Notably, SWAN achieves 2×2$\times$2 × speedups compared to Adam in terms of tokens processed when training 350M- and 1.3B-parameter models. These findings serve as a proof-of-concept that the stateless approach has the potential to serve as a practical and efficient alternative to other optimizers that require tracking internal states. Future work may explore other design choices for stateless optimizers and further expand SWAN’s applicability to other complex training regimes beyond standard LLM pre-training.
@@ Appendix A Desired properties of adaptive optimizers
Appendix B Additional Analysis
B.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
@@ Appendix B Additional Analysis
Unlike the convergence rates of GD and Adam presented in Zhang et al. (2024a), as well as Theorem 3 and Corollary 1 in Appendix, the optimal convergence rate (13) of $\mathtt{GradWhitening}$ no longer explicitly depends on the condition number $\kappa$ of $H$. In fact, consider a lower bound $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$, since trace of $H$ appear both in the nominator and denominator, we expect that to be more robust to ill-conditioned problems. For example, consider the specific initialization ${\mathbf{W}}_{\text{whitened}}^{(t)}=cI$, it is straightforward to show that $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\perp\kappa$, which is completely disentangled from the condition number. Hence $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$ would not shrink as $\kappa\xrightarrow{}\infty$. See Proposition 2 for less extreme situations.
@@
As pointed out by Zhang et al. (2024a) and Da Silva & Gazeau (2020), Adam with $\beta_{2}<1$ will have issues with convergence, which will not be completely removed even with lr decay. Therefore, we will not discuss the case of $\beta_{2}<1$ to avoid the complication.
B.2 Numerical Verification of Proposition 1
Given a STB, we consider the following standard full-batch learning dynamics (Tian et al., 2023). Define the conditional expectation $\mathbb{E}_{q=m}[\cdot]:=\mathbb{E}[\cdot|q=m]$. Consider the dynamics of the weight matrix $W$ and the attention logits $z_{q}$, if we train the model with a batch of inputs that always end up with query $q[i]=m$. The weight update for $W$ and $z_{q}$ are given by the following noisy updates:
$\dot{{\mathbf{W}}}^{(t)}=\mathbb{E}_{q=m}\left[{\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}\right],\quad\dot{{\bm{z}}}_{m}^{(t)}=\mathbb{E}_{q=m}\left[\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}\right],$ (15)
Where ${\bm{f}}^{(t)}=\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q}^{(t)})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)$, $({\bm{h}}^{(t)})^{\prime}=\phi^{\prime}(({\mathbf{W}}^{(t)})^{\top}{\bm{f}}^{(t)})$ is the derivative of the current activation, ${\mathbf{G}}_{{\bm{h}}}^{(t)}=\nabla_{{\bm{h}}^{(t)}}\mathcal{L}$ is the gradient of the loss function $\mathcal{L}$ with respect to the hidden activation ${\bm{h}}^{(t)}$, and ${\bm{g}}_{{\bm{f}}^{(t)}}^{(t)}=\sum_{k}{\bm{g}}_{{\bm{h}}_{k}^{(t)}}^{(t)}({\bm{h}}^{(t)}_{k})^{\prime}{\bm{w}}_{k}^{(t)}$ is the sum of the gradients with respect to the attention logits. Here, ${\bm{w}}_{k}^{(t)}$ is the $k$-th column of ${\mathbf{W}}^{(t)}$, ${\bm{g}}_{h_{k}^{(t)}}^{(t)}[i]$ be the backpropagated gradient sent to node $k$ at sample $i$.
Then, we numerically solving the STB ODE with $n=12,M_{C}=10$ in Equation 15. During all training steps, we analytically track the evolution of Hessian of $rmW$. Results are shown in Figure 9. As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
@@ Appendix D Proof of Theorem 2
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}}),$
noticing that $\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}=1$, and $\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})=\mathcal{P}({\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{O}})={\mathbf{Q}}{\mathbf{Q}}^{\top}{\mathbf{O}}={\mathbf{O}}$. Hence:
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-{\mathbf{O}}=\mathbf{0}={\mathbf{W}}^{*}.$
@@ Appendix F Proof of Proposition 3
Theorem 3 (Contraction factor lower bound for gradient descent, generalized based on Zhang et al. (2024a)).
@@
The proposition 1 in Zhang et al. (2024a) has shown that the lower bound holds for diagonal positive definite Hessian ${\mathbf{H}}$. To show that the lower bound holds for a general positive definite Hessian ${\mathbf{H}}$ we will reformulate the problem to align with the setup in diagonal case (Prposition 1 of Zhang et al. (2024a)).
@@
Thus, by applying Proposition 1 of Zhang et al. (2024a) to this transformed problem, we conclude that there exists initial point such that the lower bound on the contraction rate
@@
The update rule of Adam with $\beta_{1}=0$ and $\beta_{2}=1$ is given by Zhang et al. (2024a):